# sPhinX: Sample Efficient Multilingual Instruction Fine-Tuning Through N-shot Guided Prompting

### v2

@@ 4.1 Setup
Base Models: We use Mistral-7B  [^4] We specifically use the v1.0 base model from here. Jiang et al. (2023) and Phi-3-small Abdin et al. (2024) base model variants and fine-tune them on different datasets based to improve multilingual performance.
Datasets: Apart from the sPhinX dataset, we use Bactrian Li et al. (2023), MultiAlpaca Wei et al. (2023b) and Aya Singh et al. (2024) instruction datasets for comparative evaluation §3.1.
@@
  • MultiAlpaca Wei et al. (2023b) is a self instruct dataset which follows the same approach as the English-only Alpaca by translating the seed instructions to 11 languages and then using GPT-3.5.-Turbo to generate responses. We fine-tune our models on the complete set of the dataset consisting of 500k datapoints.
  • Aya Singh et al. (2024) consists of human-curated prompt-completion pairs in 65 languages called the Aya Dataset. It also consists of an aggregation of 44 monolingual and multilingual templated instruction datasets and 19 translated datasets ranging over 114 languages. The total size of the Aya dataset is around 513M instances. Due to the in-feasibility of fine-tuning our models on the entire dataset, we sampled it down to 2.7M instances, ensuring parity with our sPhinX dataset by selecting equal numbers of samples for each language in our subset.
Evaluation: We evaluate [^5] Evaluation prompts and other details in Appendix §A.1 our fine-tuned models along with the available Instruction fine-tuned model variants of Mistral-7B and Phi-3-small (IFT [^6] We take the Mistral-7B instruction-tuned variant from here and Phi-3-small variant from here.) on XCOPA Ponti et al. (2020), XStoryCloze Lin et al. (2022), XWinograd Muennighoff et al. (2023)Tikhonov and Ryabinin (2021), Belebele Bandarkar et al. (2023), and XQuAD Artetxe et al. (2020) using the language model evaluation harness Gao et al. (2023).
@@
4.2 Fine-Tuning Methodology
Inspired from  Longpre et al. (2023)’s strategies to instruction tune a model, we devise Language-Specific N-shot Guided Instruction fine-tuning (LANG). With this approach, we augment a training example by prepending $N$ number of samples of same language as that of the original training example randomly selected from the corresponding few shot set. This augmented training example is used for instruction tuning our models. Suppose a training example in language $l$ is a pair of $Instruction({I}_{\text{train}}^{l})$ and $Response({R}_{\text{train}}^{l})$. We need to prepend $N$ number of shots represented by $({I}_{\text{fewshot}_{1}}^{l},{R}_{\text{fewshot}_{1}}^{l})$, $({I}_{\text{fewshot}_{2}}^{l},{R}_{\text{fewshot}_{2}}^{l})$, …, $({I}_{\text{fewshot}_{N}}^{l},{R}_{\text{fewshot}_{N}}^{l})$. For every training example, $N$ is chosen with a probability $p$ as defined in Table 3. $\mathcal{T}$ is the instruction tuning templating function which takes $Instruction$ and $Response$ and transforms it to user-assistant format by adding the special tokens. Thus, the final training example is:
$(\bigoplus_{i=1}^{N}\mathcal{T}(I_{\text{fewshot}_{i}}^{l},R_{\text{fewshot}_{i}}^{l})\oplus\mathcal{T}(I_{\text{train}}^{l},R_{\text{train}}^{l}))$
where $N$ is chosen based on the probability distribution $P(N)$. $\bigoplus$ represents combining all few-shot examples, while $\oplus$ indicates concatenating the main training example with the aggregated few-shot examples. Once $N$ is selected, an equivalent number of few-shot examples is sampled uniformly at random from the few shot set. The maximum context token length for training the models is set at 8192. To ensure augmented samples fall within this range, we assign more weights to $N$=0 and $N$=1. This strategy is consistently applied when instruction tuning other datasets for both Phi-3-small and Mistral-7B base model. The average token length of samples in each datasets as per Phi-3-small model tokenizer is shown in Table 1. To assess the effectiveness of each instruction tuning dataset on an equal scale, we conducted a comparative analysis of model performance, fine-tuning each on approximately 8 billion tokens from each dataset using LANG strategy.
4.3 Hyperparameters and Training Setup
@@
5 Results
Figure 4: Performance of Mistral-7B and Phi-3-smallwhen instruction-tuned on 8B tokens across various datasets on different benchmarks.
Figure 5: Overall performance for both Mistral-7B and Phi-3-small on our dataset and fine-tuning recipe. Legend: Base_Model: Base Model, IFT: Instruction Fine-Tuned variant of the Base Model available publicly, Bactrian: Base model trained on Bactrian dataset, mAlpaca: Base Model trained on mAlpaca dataset, Aya: Base Model trained on a subset of the Aya dataset, Sphinx-0s: Base Model trained in a 0-shot fashion on the sPhinX dataset, Sphinx: Base model trained using LANG on sPhinX dataset
We evaluate the reasoning, question answering, and reading comprehension abilities of the Phi-3-small and Mistral-7B models, instruction-tuned on different multilingual datasets, using various benchmarks and find that fine-tuning on sPhinX provides an improvement of 4.2%pt [^8] A percentage point (pt) is the unit for expressing the absolute difference between two percentage values and 5%pt respectively ( language-wise results are in the Appendix §A.2). Additionally as observed in Figure 4, the sPhinX dataset significantly enhances the multilingual performance of the Phi-3-small and Mistral-7B model compared to other datasets even when finetuned on equal number of tokens. Furthermore, during instruction tuning on 8B tokens, the models encountered fewer training samples for sPhinX due to its higher average token length per sample, as illustrated in Table 1.
Due to the code-mixed nature of the instruction along with CoT reasoning explanations, a single sample of sPhinX is notably richer as compared to its counterparts from the other datasets. Consequently, even with fewer samples (keeping the number of the tokens same), models trained on sPhinX achieve better performance; thereby demonstrating the per sample efficiency of sPhinX.
6 Ablation: Improvements from LANG
To demonstrate the effectiveness of our LANG strategy, we also instruction-tuned the models on sPhinX without any pre-pended shots, referring to this as sPhinX-0s. As shown in Figure 5 (with detailed results in Appendix §A.2), models fine-tuned on sPhinX especially Mistral-7B exhibits superior performance compared to its counterparts fine-tuned on other datasets across all benchmarks. Moreover, fine-tuning both Phi-3-small and the Mistral-7B on sPhinX using the LANG strategy further boosts the performance by 3.2%pt and 10%pt respectively as compared to the base model fine-tuned without this strategy. Surprisingly, even without the LANG strategy, models instruction-tuned on sPhinX still perform better than those tuned on other datasets with LANG.
Furthermore, employing the LANG strategy leads to additional performance improvements indicating that LANG can effectively enhance the multilingual capabilities of LLMs. From the detailed results in Appendix §A.2, we observe no performance regression on high resource languages which normally occurs due to catastrophic forgetting Chang et al. (2023).
We also observe significant performance improvements in medium and low-resource languages such as Arabic, Hindi, Thai, Turkish, Tamil, and Telugu, further showcasing the effectiveness of our dataset and the LANG fine-tuning strategy (Appendix §A.2).
6.1 Regression Analysis on Standard LLM Benchmarks
It is well-studied that training on multiple languages cause regression on the performance on English language due to catastrophic forgetting Chang et al. (2023). We test this by checking for performance of the Phi-3-small model fine-tuned with sPhinX on English in the multilingual benchmarks we evaluate ((Appendix §A.2) and on popular English-only benchmarks.
We find that the Phi-3-small sPhinX model maintains its performance in English on the multilingual benchmarks and is also consistently able to maintain performance on standard English benchmarks such as MMLU (5-shot) Hendrycks et al. (2021), MedQA (2-shot) Jin et al. (2021), Arc-C (10-shot), Arc-E (10-shot) Clark et al. (2018), PiQA (5-shot) Bisk et al. (2020), WinoGrande (5-shot) Sakaguchi et al. (2021), OpenBookQA (10-shot) Mihaylov et al. (2018), BoolQ (2-shot) Clark et al. (2019) and CommonSenseQA (10-shot) Talmor et al. (2018) (Table 10). We notice some regression in the GSM-8k (8-shot, CoT) Cobbe et al. (2021) benchmark. This indicates that gains in multilingual performance caused by sPhinX do not come at the cost of regression in English performance.
7 Conclusion
In this paper, we demonstrated how instruction tuning Phi-3-small and Mistral-7B on sPhinX effectively improve their multilingual capabilities. We observed that instruction tuning the models using the sPhinX dataset leads to consistent gains on an average by 4.2%pt for Phi-3-small and 5%pt for Mistral-7B respectively. We find that the Phi-3-small model instruction tuned on sPhinX is the best performing model on multilingual benchmarks. Moreover, sPhinX exhibits greater sample efficiency and diversity compared to other multilingual instruction tuning datasets.
Additionally, we proposed a method for further enhancing the model’s performance by utilizing LANG, which supplements the training examples with $N$ samples from a few-shot set, providing the model with additional context to aid in its learning process. This further boosts the performance for both Phi-3-small and Mistral-7B by 3.2%pt and 10%pt respectively. Models instruction-tuned on sPhinX, exhibit enhanced performance even in languages they have not previously encountered during training. Finally, we evaluate the performance of the Phi-3-small model fine-tuned on sPhinX and find that it is able to maintain English performance, suggesting that gains in multilingual performance do not come at the cost of English performance while fine-tuning with sPhinX. Through our comprehensive experiments, analyses, and findings, we aim to contribute to the progress of LLMs for multilingual purposes, promoting advancements in natural language processing across a broader spectrum of languages.
8 Future Work
We have conducted all experiments using 7B base models in full parameter fine-tuning settings. It would be interesting to study the same effect by adaptive learning using LoRA Hu et al. (2022) or PEFT Mangrulkar et al. (2022). We believe that our strategies could also be effective for models with fewer parameters, leading to a notable improvement in multilingual performance.
Our LANG strategy involves utilizing $N$ examples from the same language. Other ideas that can be explored are incorporating $N$ examples from the same language script to increase the diversity of the sample. This approach could be particularly beneficial for enhancing the performance of models in low-resource languages.
Limitations
Our study has several limitations that can be considered in future research. Firstly, we conducted an extensive series of experiments, utilizing significant GPU resources and substantial time for model fine-tuning. Due to these resource-intensive processes, it may be difficult to apply our strategies to fully fine-tune a model. Besides, our study is confined to 7B models, explicitly excluding larger models. Despite this limitation, we believe our methodologies are broadly applicable for fine-tuning smaller datasets using techniques like LoRA and PEFT. Secondly, our fine-tuning dataset focuses on reasoning tasks and excludes some low-resource languages. We evaluated the models’ performance against these reasoning benchmarks. However, we did not benchmark our models on generative tasks such as summarization, nor did we evaluate models on hallucination, toxicity, or fairness.
Ethics Statement
Despite our rigorous efforts to ensure that our dataset is free from discriminatory, biased, or false information, there remains a possibility that these problems are present, particularly in multilingual contexts. Hence, it is possible that these issues might propagate to our fine-tuned models as well. We are committed to mitigate such risks and strongly advocate for the responsible use of recipes and prevent any unintended negative consequences.
Acknowledgements
We thank the Orca team at Microsoft Research for sharing the Orca dataset to implement the synthetic dataset creation recipe.
Appendix A Appendix
A.1 Prompt Templates
Figure 6 is the template for Selective Translation that was used to generate the synthetic data. Our reference dataset is in English and the {language} is the target language to generate the data in. Figure 7, 8, 9,10 and 11 are the prompts used to evaluate XQuAD, XstoryCloze, Xwinograd, XCOPA and Belebele respectively.
A.2 Detailed Results
Tables 4, 5, 6, 7 and 8 show the granular results on our models and dataset.

### v4

@@ 5.1 Setup
Base Models: We use Mistral-7B [^3] We specifically use the v1.0 base model from
https://huggingface.co/mistralai/Mistral-7B-v0.1
and Phi-3-small Abdin et al. (2024) base model variants and instruction fine-tune them.
Datasets: Apart from the sPhinX dataset, we use Bactrian Li et al. (2023), m-Alpaca Wei et al. (2023b) and Aya Singh et al. (2024b) instruction datasets for comparative evaluation. We also utilize the Azure Translator API [^4] https://azure.microsoft.com/en-us/products/ai-services/ai-translator (sPhinX-T) to translate the original dataset into all our target languages, demonstrating the effectiveness of our Selective Translated Augmentation approach. More details about the datasets used for comparative evaluation are present in Appendix §A.2.
Evaluation: We evaluate [^5] Evaluation prompts and other details in Appendix §A.1 and §A.3. our fine-tuned models along with the available base and Instruction fine-tuned model variants of Mistral-7B and Phi-3-small (IFT [^6] We take the Mistral-7B instruction-tuned variant from https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1 and Phi-3-small variant from https://huggingface.co/microsoft/Phi-3-small-8k-instruct.) on 4 discriminative tasks; XCOPA Ponti et al. (2020)(4-shot), XStoryCloze Lin et al. (2022)(4-shot), XWinograd Muennighoff et al. (2023)(0-shot), Tikhonov and Ryabinin (2021), Belebele(0-shot) Bandarkar et al. (2023), and 2 generative tasks; XQuAD (3-shot) Artetxe et al. (2020) and Translation (4-shot) Bojar et al. (2014, 2016); Kocmi et al. (2023) using the language model evaluation harness Gao et al. (2023). The number of few-shot selection are inspired from these works Ahuja et al. (2023, 2024); Asai et al. (2024)
Apart from generative tasks such as XQuAD, and machine translation, we also evaluate our instruction-tuned models on open-ended generation prompts. For this, we use an LLM-based evaluation approach to simulate win rates. We use the open-source test set from the Aya Dataset (Singh et al., 2024a), which includes 250 prompts per language across six languages. We use GPT-4o as the LLM evaluator to pick the preferred model generation on this test set, and we subsequently compute win rates (%) based on these preferences. To avoid a potential bias, we randomize the order of the models during the evaluation. The prompt for the evaluator is described in the Appendix: §A.1.
6 Results
Figure 3: Performance of Mistral-7B and Phi-3-small when instruction-tuned on 8B tokens across various datasets on different benchmarks. The solid lines represent the Phi fine-tuned models and the dashed lines represent the Mistral fine-tuned models.
Model $XC$ $XS$ $XW$ $XQ$ $BL$ $MT^{1}$ $MT^{2}$
Mistral-7B
Base Model 0.63 0.68 0.52 0.74 0.24 0.54 0.42
IFT 0.62 0.73 0.54 0.60 0.47 0.49 0.39
m-Alpaca 0.55 0.59 0.51 0.46 0.41 0.41 0.39
Aya 0.68 0.71 0.54 0.66 0.38 0.39 0.37
Bactrian 0.54 0.67 0.54 0.69 0.26 0.45 0.34
sPhinX-T 0.61 0.78 0.57 0.78 0.67 0.49 0.38
sPhinX-0s 0.58 0.58 0.68 0.69 0.67 0.49 0.42
sPhinX 0.68 0.81 0.71 0.80 0.71 0.55 0.46
Phi-3-small
Base Model 0.64 0.78 0.75 0.78 0.65 0.54 0.42
IFT 0.68 0.79 0.78 0.75 0.70 0.54 0.46
m-Alpaca 0.68 0.79 0.81 0.77 0.75 0.45 0.39
Aya 0.65 0.79 0.69 0.83 0.72 0.41 0.40
Bactrian 0.71 0.82 0.73 0.85 0.77 0.54 0.40
sPhinX-T 0.70 0.80 0.77 0.78 0.75 0.55 0.44
sPhinX-0s 0.71 0.81 0.80 0.82 0.79 0.56 0.45
sPhinX 0.72 0.84 0.87 0.87 0.79 0.56 0.46
Table 1: Performance of Mistral-7B and Phi-3-small instruction-tuned on various datasets. Abbreviations: $XC$ XCOPA (Acc.,4-shot), $XS$ XStoryCloze (Acc.,4-shot) , $XW$ XWinograd (Acc., 0-shot), $XQ$ XQuAD (F1,3-shot), $BL$ Belebele (Acc., 0-shot). $MT^{1}$- Translation for x:en (ChrF, 4-shot), $MT^{2}$ Translation for en:x direction (ChrF, 4-shot). The best performing dataset for each model is indicated in bold, and the overall best performing model is indicated with an underline.
We evaluate reasoning, question answering, translation and reading comprehension abilities of the Phi-3-small and Mistral-7B models, instruction-tuned on different multilingual datasets, using various benchmarks and find that fine-tuning on sPhinX provides an average improvement of 39.8% and 11.2% respectively on both the models. (Refer to sPhinX-0s in Table 1 for overall results and to Appendix §A.5 for language-wise results). Additionally, as observed in Figure 3, the sPhinX dataset significantly enhances the multilingual performance of the Phi-3-small and Mistral-7B model compared to other datasets even when fine-tuned on an equal number of tokens.
7 Ablations
7.1 Improvements from LANGIT
To demonstrate the effectiveness of our LANGIT strategy, we also instruction-tuned the models on sPhinX with 0 shots, referring to this as sPhinX-0s. As shown in Table 1 (with detailed results in Appendix §A.5), models fine-tuned on sPhinX especially Mistral-7B exhibit superior performance compared to its counterparts fine-tuned on other datasets across all benchmarks. Moreover, fine-tuning both Mistral-7B and the Phi-3-small on sPhinX using the LANGIT strategy further boosts the performance by an average of 15% and 3.2% respectively as compared to the vanilla fine-tuned model (sPhinX -0s) across multilingual benchmarks.
Furthermore, employing the LANGIT strategy leads to additional performance improvements indicating that LANGIT can effectively enhance the multilingual capabilities of LLMs. From the detailed results in Appendix §A.5, we observe no performance regression on high resource languages which normally occurs due to catastrophic forgetting Chang et al. (2023).
We also observe significant performance improvements in medium and low-resource languages such as Arabic, Hindi, Thai, Turkish, Tamil, and Telugu, further showcasing the effectiveness of our dataset and the LANGIT fine-tuning strategy (Appendix §A.5).
7.2 Comparisons with the API translated dataset
Due to the code-mixed nature of the instruction along with CoT reasoning explanations, a single sample of sPhinX is notably richer as compared to its counterparts from the other datasets. This can be observed in the Table 1 for sPhinX-T wherein the sPhinX trained models with the LANGIT strategy outperform the directly translated dataset baselines by an average of 11.7% and 6.3% for both Mistral-7B and Phi-3-small respectively when compared to the sPhinX-T baselines across multilingual benchmarks. Consequently, even with fewer samples as compared to the other datasets (keeping the number of the tokens the same), models trained on sPhinX achieve better performance, thereby demonstrating the per-sample efficiency of sPhinX.
7.3 Simulated Preference Evaluation
As shown in Table 2, our win-rate experiments reveal that the GPT-4o evaluator predominantly favored outputs generated by the Mistral base model trained on the sPhinX dataset using the LANGIT strategy over other models. For the Phi baselines, we observed a higher percentage of TIEs for all the languages except English, where the evaluator rated both outputs equally, rather than favoring a specific model. This performance gap between the Mistral and Phi models likely arises from the age of their respective base models. Since the Mistral base model is older, it benefits more from additional training on our dataset, whereas the more recently released Phi models are already competitive enough on these benchmarks resulting in preferring both the outputs equally.
Model ar en po te tu zh
Mistral-7B
IFT 75 / 9 / 16 59 / 36 / 5 57 / 27 / 16 62 / 15 / 23 65 / 13 / 22 70 / 26 / 4
m-Alpaca 85 / 2 / 13 74 / 21 / 5 52 / 33 / 15 69 / 8 / 23 70 / 4 / 25 75 / 15 / 10
Aya 78 / 11 / 11 85 / 11 / 4 55 / 30 / 15 56 / 18 / 25 62 / 19 / 19 78 / 14 / 8
Bactrian 82 / 4 / 13 85 / 12 / 3 57 / 31 / 12 56 / 18 / 26 62 / 20 / 18 74 / 14 / 12
sPhinX-T 61 / 23 / 16 63 / 27 / 10 56 / 24 / 20 56 / 24 / 20 62 / 19 / 19 66 / 23 / 11
sPhinX-0s 65 / 19 / 16 74 / 20 / 5 55 / 31 / 14 51 / 22 / 27 52 / 36 / 12 68 / 20 / 11
Phi-3-small
IFT 46 / 38 / 17 55 / 43 / 2 50 / 44 / 6 39 / 29 / 36 30 / 27 / 43 50 / 44 / 6
m-Alpaca 46 / 5 / 49 59 / 29 / 12 44 / 21 / 35 33 / 6 / 60 31 / 10 / 59 30 / 6 / 64
Aya 40 / 18 / 42 80 / 11 / 9 56 / 16 / 28 37 / 18 / 46 25 / 21 / 54 36 / 22 / 41
Bactrian 32 / 13 / 55 68 / 17 / 15 51 / 14 / 36 24 / 14 / 62 32 / 15 / 53 26 / 11 / 63
sPhinX-T 35 / 14 / 51 75 / 12 / 13 61 / 12 / 27 23 / 14 / 63 36 / 18 / 47 37 / 12 / 51
sPhinX-0s 23 / 11 / 66 61 / 17 / 22 32 / 18 / 50 14 / 9 / 77 15 / 10 / 74 14 / 7 / 79
Table 2: Win rates (%) according to GPT-4o: The first value represents the percentage of outputs where the evaluator preferred the sPhinX and LANGIT trained model. The second value indicates the percentage of outputs preferred from the target model. The third value reflects cases where the evaluator rated both outputs equally (TIE).
Benchmarks Base Model sPhinX
MMLU
(5-shot)
0.76 0.75
HellaSwag
(5-shot)
0.81 0.83
GSM-8k
(8-shot, CoT)
0.85 0.77
MedQA
(2-shot)
0.64 0.66
Arc-C
(10-shot)
0.90 0.90
Arc-E
(10-shot)
0.97 0.97
PIQA
(5-shot)
0.84 0.89
WinoGrande
(5-shot)
0.77 0.82
OpenBookQA
(10-shot)
0.86 0.88
BoolQ
(2-shot)
0.82 0.87
CommonSenseQA
(10-shot)
0.80 0.81
7.4 Regression Analysis on Standard LLM Benchmarks
It is well studied that training in multiple languages causes regression in performance in English due to catastrophic forgetting Chang et al. (2023). We test this phenomenon for our trained models by checking the performance of the Phi-3-small model fine-tuned with sPhinX on English in the multilingual benchmarks we evaluate ((Appendix §A.5) and on popular English-only benchmarks (Table 3).
We find that the Phi-3-small fine-tuned on sPhinX maintains its performance in English on the multilingual benchmarks and is also consistently able to maintain performance on standard English benchmarks such as MMLU (5-shot) Hendrycks et al. (2021), MedQA (2-shot) Jin et al. (2021), Arc-C (10-shot), Arc-E (10-shot) Clark et al. (2018), PiQA (5-shot) Bisk et al. (2020), WinoGrande (5-shot) Sakaguchi et al. (2021), OpenBookQA (10-shot) Mihaylov et al. (2018), BoolQ (2-shot) Clark et al. (2019) and CommonSenseQA (10-shot) Talmor et al. (2018) (Table 3). We notice some regression in the GSM-8k (8-shot, CoT) Cobbe et al. (2021) benchmark. This indicates that gains in multilingual performance caused by sPhinX do not come at the cost of regression in English performance.
8 Conclusion
In this paper, we demonstrated how instruction tuning Mistral-7B and Phi-3-small on sPhinX effectively improve their multilingual capabilities. We observed that instruction tuning the models using the sPhinX dataset leads to performance improvement by an average of 39.8% and 11.2% for Mistral-7B and Phi-3-small respectively when compared to their corresponding base models across multilingual benchmarks. Moreover, sPhinX exhibits greater sample efficiency and diversity compared to other multilingual instruction tuning datasets. We also proposed LANGIT, a strategy that enhances model performance by incorporating $N$ few-shot examples, boosting results by 15% and 3.2% for Mistral-7B and Phi-3-small, respectively, over vanilla fine-tuning with sPhinX. Compared to the sPhinX-T translation baseline, LANGIT yielded gains of 11.7% and 6.3%. Models fine-tuned on sPhinX also showed improved performance in unseen languages without degrading English performance. We also observed that the GPT-4o win-rate evaluations favored Mistral-7B with LANGIT, while Phi-3-small showed more ties due to its stronger baseline. Finally, we plan on releasing a subset of our augmented dataset built on OpenOrca (Open-sPhinX).
9 Future Work
All experiments were conducted using 7B base models with full fine-tuning. It would be interesting to explore our methods with adaptive fine-tuning techniques like LoRA Hu et al. (2022) or PEFT Mangrulkar et al. (2022), and on smaller models, where we expect similar gains in multilingual performance. Our LANGIT strategy uses $N$ examples from the same language and future work could investigate using $N$ examples from the same script to introduce greater diversity, especially for improving performance in low-resource languages. We also observed code-switched data being generated when we employed this strategy to generate data. It will be interesting to explore this phenomena in a future study.
Limitations
Our study has several limitations that can be considered in future research. Firstly, we conducted an extensive series of experiments, utilizing significant GPU resources and substantial time for model fine-tuning. Due to these resource-intensive processes, it may be difficult to apply our strategies to fully fine-tune a model. Besides, our study is confined to 7B models, explicitly excluding larger models. Despite this limitation, we believe that our methodologies are broadly applicable for fine-tuning smaller datasets using techniques like LoRA and PEFT. While some languages were not included, we made a conscious effort to cover a diverse set spanning multiple scripts and language families.
Ethics Statement
Despite our rigorous efforts to ensure that our dataset is free from discriminatory, biased, or false information, there remains a possibility that these problems are present, particularly in multilingual contexts. Hence, it is possible that these issues might propagate to our fine-tuned models as well. We are committed to mitigating such risks and strongly advocate for the responsible use of recipes and prevent any unintended negative consequences.
Appendix A Appendix
A.1 Prompt Templates
Figure 4 is the template for Selective Translated Augmentation that was used to generate the synthetic data. Our reference dataset is in English and the {language} is the target language to generate the data in. Figure 5, 6, 7,8, 9 and 10 are the prompts used to evaluate XQuAD, XstoryCloze, Xwinograd, XCOPA, Belebele and Translation respectively. Figure 11 denotes the prompt used for simulating win-rate evaluations.
A.2 Baseline Datasets For Comparative Evaluation
@@
  • m-Alpaca Wei et al. (2023b) is a self-instruct dataset that translates seed instructions from English to 11 languages, using GPT-3.5-Turbo for response generation. We fine-tune our models on the full dataset, which contains 500k data points.
  • Aya Singh et al. (2024a) contains human-curated prompt-completion pairs in 65 languages, along with 44 monolingual and multilingual instruction datasets and 19 translated datasets across 114 languages, totaling around 513M instances. To ensure parity with the sPhinX dataset, we sampled it down to 2.7M instances, ensuring equal representation for each language in our subset.
A.3 Evaluation Benchmarks
@@
  • Translation: We utilize a subset of WMT14, WMT16 and WMT23 of language pairs (7 languages), with evaluation in a 4-shot setting.
A.4 Hyperparameters and Training Setup
@@
A.5 Detailed Results
Tables 8, 9, 10, 11, 12, 13 and 14 show the granular results on our models and dataset.